OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r = <8, 1, -6>; v = <6, 7, -3>; w = <-7, 5, 2>

jimthompson5910 (jim_thompson5910):

if x = <a,b,c> y = <d,e,f> then x dot y = a*d + b*e + c*f

jimthompson5910 (jim_thompson5910):

essentially you multiply the corresponding elements, then add up the products to get the final answer

OpenStudy (anonymous):

So would it be 8*6+1*7+-6*-3=

jimthompson5910 (jim_thompson5910):

that's if you wanted to find r dot v however you want to find v dot w

OpenStudy (anonymous):

Oh so it's 6*-7+7*5+-3*2=-13

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome